OpenStudy (anonymous):

Evaluate. 90 • 4 – 14 ÷ 2 + 102 ÷ 5 • 2

OpenStudy (anonymous):

help! me

OpenStudy (anonymous):

first do multiplying and dividing, then sum and minus :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i need more help

OpenStudy (anonymous):

help

OpenStudy (anonymous):

90*4 = 360 14/2 =7 102/10 = 10,2 360 - 7 + 10.2 = ...

OpenStudy (anonymous):

can you do the rest by yourself? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, post the answer here so i can tell you if it's correct :)

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

this problem challenges your understanding of "order of operations" rules. Could you look up that topic? some people refer to these rules with the acronym "PENDAS."

OpenStudy (anonymous):

363.2

OpenStudy (anonymous):

well done, it's correct :)

OpenStudy (anonymous):

yay thanks

OpenStudy (anonymous):

no problem :)